# Volume polynomials

### v1

@@ 4. Linear operators preserving volume polynomials
where $\alpha\leq\beta$ means that their components satisfy $\alpha_{i}\leq\beta_{i}$ for all $i\in E$. For any further conventions for multivariate polynomials, we refer to [BH20, Section 2]. For $\mu\in\mathbb{Z}^{n}_{\geq 0}$, we consider
@@
As observed in [Alu24, Remark 2.2], the property of being a realizable covolume polynomial over $k$ does not depend on the choice of $\mu$. This follows from the translation invariance
@@
For the cone construction that justifies this, see [GHM+, Section 2].
@@
According to [BKH+, Section 2.1], the dual of an $\mathrm{M}$-convex set $J\subseteq\mathbb{Z}^{E}_{\geq 0}$, defined up to translation in $\mathbb{Z}^{E}_{\geq 0}$, is the $\mathrm{M}$-convex subset
@@
where $\mu$ is any nonnegative integral vector satisfying $\alpha\leq\mu$ for all $\alpha\in J$. [^6] A standard choice used in [BKH+, Section 2.1] is to take $\mu$ to be the duality vector $\delta_{J}=\delta^{+}_{J}+\delta^{-}_{J}$, where $\delta^{+}_{J}=\sup J$ and $\delta^{-}_{J}=\inf J$. Since the support of a covolume polynomial over $k$ is the dual of the support of a volume polynomial over $k$, the support of a covolume polynomial is an $\mathrm{M}$-convex set.
@@
The class of (realizable) volume polynomials over $k$ is closed under nonnegative (rational) linear changes of coordinates, as well as under taking products and minors. Similarly, the class of (realizable) covolume polynomials over $k$ is closed under nonnegative (rational) linear changes of coordinates [GHM+, Theorem 2.7], as well as taking products and minors [GHM+, Theorem 1.5]. It is interesting to note that the corresponding statements for volume polynomials and covolume polynomials sometimes have substantially different proofs.
@@
Let $f_{J}$ be the basis generating polynomial of a matroid $J$ linear over $k$. By Example 3.5, $f_{J}$ is a realizable volume polynomial over $k$. Since the dual of a linear matroid over $k$ is linear over the same field [Oxl11, Corollary 2.2.9], $f_{J}$ is a realizable covolume polynomial over $k$ as well.
@@
The proof of [HMMSD22, Theorem 6] shows that the Schubert polynomial $s_{w}(\partial)$ is a realizable covolume polynomial over $k$ for any permutation $w$ and any $k$. In particular, any Schur polynomial is a realizable covolume polynomial over $k$ for any field $k$. By Example 3.6, any normalized Schur polynomial is a realizable volume polynomial over $k$ for any $k$. It is not known whether normalized Schubert polynomials are realizable volume polynomials over $k$ for any $k$. See [HMMSD22, Conjecture 15] for a weaker statement.
@@
The main result of [GHM+] is the following characterization of realizable covolume polynomials. This parallels the characterization of dually Lorentzian polynomials in [RSW25, Theorem 1.2].
@@
Example 3.11 shows that the set of volume polynomials of convex bodies does not satisfy the analogous statement. [^7] For example, $\partial_{5}f_{C}$ is not a volume polynomial of convex bodies. For a parallel statement characterizing volume polynomials as linear operators preserving covolume polynomials, see [GHM+, Theorem 1.9].
Corollary 4.8 can be used to deduce new inequalities for mixed volumes of convex bodies, or more generally, for intersection numbers of nef divisors on a projective variety. For instance, given a Schubert polynomial $s_{w}(\partial)$ and a volume polynomial $f_{C}(x)$, any known inequality for the coefficients of a volume polynomial can be applied to $s_{w}(\partial)\circ f_{C}(x)$ to produce another inequality for the coefficients of $f_{C}(x)$. For an overview of known inequalities for the coefficients of the volume polynomial, such as the Khovanskii–Teissier inequality or the reverse Khovanskii–Teissier inequality, see [HMWX].
@@
The symbol theorem for Lorentzian polynomials states that, if the symbol of a linear operator $T$ is a Lorentzian polynomial, then $T$ sends Lorentzian polynomials to Lorentzian polynomials [BH20, Theorem 3.2]. Theorem 4.7 can be used to derive a volume polynomial analogue of the symbol theorem for Lorentzian polynomials. [^8] The study of symbols of linear operators dates back to Gårding [Gar51] and appears prominently in the work of Borcea and Brändén on the Pólya–Schur program for stable polynomials [BB09a, BB09b].
@@
then, by [GHM+, Lemma 2.1], it preserves the classes of irreducible cycles up to a rational multiple.
@@
    The upper truncation operators and the lower truncation operators preserve realizable volume polynomials over $k$ [GHM+, Corollary 3.3].
@@
    The polarization operator $\Pi^{\uparrow}$ preserves realizable volume polynomials over $k$ [GHM+, Proposition 4.1].
@@
    The normalization operator $\mathrm{N}$ preserves realizable volume polynomials over $k$ [GHM+, Proposition 4.2].
@@
    For any nonnegative rational number $t$, the interlacing operator $1+tx_{i}\partial_{j}$ preserves realizable volume polynomials over $k$ [GHM+, Proposition 4.3].
@@
    For any nonnegative rational number $t$, the symmetric exclusion process $\Phi^{i,j}_{t}$ preserves realizable multiaffine volume polynomials over $k$ [GHM+, Proposition 4.4].
@@ 5. Realization problems for polymatroids
The following connection between algebraic matroids and the support of realizable volume polynomials is known [GHM+, Proposition 5.4]. A polymatroid on $E$ is algebraic over $k$ if there are field extensions $k\subseteq\ell_{i}\subseteq\ell$ for $i\in E$ such that
@@
Proposition 5.5 implies that every minor of an algebraic polymatroid over $k$ is an algebraic polymatroid over $k$ [GHM+, Section 5]. [^10] Definition 3.9, applied to the basis generating polynomials of polymatroids, define the notion of minor of polymatroids [BKH+, Section 2.1]. In the classical case of matroids, this statement is typically deduced from a theorem of Lindström [Lin89], who proved Piff’s conjecture that $M$ is algebraic over $k$ if $M$ is algebraic over an extension of $k$, see [Oxl11, Corollary 6.7.14]. Since the set of realizable volume polynomials over $k$ depends only on the characteristic of $k$ [GHM+, Proposition 2.10], Proposition 5.5 gives the following version of Lindström’s theorem for polymatroids.
@@
Another consequence of Proposition 5.5 is that the intersection $M_{1}\wedge M_{2}$ of algebraic matroids over $k$ is an algebraic matroid over $k$ [GHM+, Theorem 5.11]. This generalizes Piff’s theorem that the truncation of an algebraic matroid is algebraic [Wel76, Section 11.3]. [^11] In [Wel76, Section 11.3], Welsh proves the dual statement that the union $M_{1}\vee M_{2}$ of algebraic matroids over $k$ is an algebraic matroid over $k$. It is interesting to note that, as in Remark 4.3, the proof of the statement for $M_{1}\wedge M_{2}$ is substantially different from that for $M_{1}\vee M_{2}$.
Is the support of a realizable covolume polynomial over $k$ an algebraic polymatroid over $k$? This question extends the following long-standing open problem in matroid theory. For up-to-date discussions, see [BFP25, Hoc].

### v2

@@ 4. Linear operators preserving volume polynomials
where $\alpha\leq\beta$ means that their components satisfy $\alpha_{i}\leq\beta_{i}$ for all $i\in E$. For any further conventions for multivariate polynomials, we refer to [9, Section 2]. For $\mu\in\mathbb{Z}^{n}_{\geq 0}$, we consider
@@
As observed in [1, Remark 2.2], the property of being a realizable covolume polynomial over $k$ does not depend on the choice of $\mu$. This follows from the translation invariance
@@
For the cone construction that justifies this, see [16, Section 2].
@@
According to [2, Section 2.1], the dual of an $\mathrm{M}$-convex set $J\subseteq\mathbb{Z}^{E}_{\geq 0}$, defined up to translation in $\mathbb{Z}^{E}_{\geq 0}$, is the $\mathrm{M}$-convex subset
@@
where $\mu$ is any nonnegative integral vector satisfying $\alpha\leq\mu$ for all $\alpha\in J$. [^6] A standard choice used in [2, Section 2.1] is to take $\mu$ to be the duality vector $\delta_{J}=\delta^{+}_{J}+\delta^{-}_{J}$, where $\delta^{+}_{J}=\sup J$ and $\delta^{-}_{J}=\inf J$. Since the support of a covolume polynomial over $k$ is the dual of the support of a volume polynomial over $k$, the support of a covolume polynomial is an $\mathrm{M}$-convex set.
@@
The class of (realizable) volume polynomials over $k$ is closed under nonnegative (rational) linear changes of coordinates, as well as under taking products and minors. Similarly, the class of (realizable) covolume polynomials over $k$ is closed under nonnegative (rational) linear changes of coordinates [16, Theorem 2.7], as well as taking products and minors [16, Theorem 1.5]. It is interesting to note that the corresponding statements for volume polynomials and covolume polynomials sometimes have substantially different proofs.
@@
Let $f_{J}$ be the basis generating polynomial of a matroid $J$ linear over $k$. By Example 3.5, $f_{J}$ is a realizable volume polynomial over $k$. Since the dual of a linear matroid over $k$ is linear over the same field [32, Corollary 2.2.9], $f_{J}$ is a realizable covolume polynomial over $k$ as well.
@@
The proof of [21, Theorem 6] shows that the Schubert polynomial $s_{w}(\partial)$ is a realizable covolume polynomial over $k$ for any permutation $w$ and any $k$. In particular, any Schur polynomial is a realizable covolume polynomial over $k$ for any field $k$. By Example 3.6, any normalized Schur polynomial is a realizable volume polynomial over $k$ for any $k$. It is not known whether normalized Schubert polynomials are realizable volume polynomials over $k$ for any $k$. See [21, Conjecture 15] for a weaker statement.
@@
Conjecture 1.11 suggests that a quadratic polynomial $\sum_{1\leq i<j\leq n}q_{ij}\partial_{i}\partial_{j}$ is a covolume polynomial over $k$ if and only if its coefficients are nonnegative and satisfy the Plücker relations over the triangular hyperfield $\mathbb{T}_{2}$:
$\sqrt{q_{ij}q_{kl}}\leq\sqrt{q_{ik}q_{jl}}+\sqrt{q_{il}q_{jk}}\ \ \text{for any $1\leq i<j<k<l\leq n$.}$
For general discussions of Grassmannians over triangular hyperfields, see [3, 4].
@@
The main result of [16] is the following characterization of realizable covolume polynomials. This parallels the characterization of dually Lorentzian polynomials in [34, Theorem 1.2].
@@
Example 3.11 shows that the set of volume polynomials of convex bodies does not satisfy the analogous statement. [^7] For example, $\partial_{5}f_{C}$ is not a volume polynomial of convex bodies. For a parallel statement characterizing volume polynomials as linear operators preserving covolume polynomials, see [16, Theorem 1.9].
Corollary 4.8 can be used to deduce new inequalities for mixed volumes of convex bodies, or more generally, for intersection numbers of nef divisors on a projective variety. For instance, given a Schubert polynomial $s_{w}(\partial)$ and a volume polynomial $f_{C}(x)$, any known inequality for the coefficients of a volume polynomial can be applied to $s_{w}(\partial)\circ f_{C}(x)$ to produce another inequality for the coefficients of $f_{C}(x)$. For an overview of known inequalities for the coefficients of the volume polynomial, such as the Khovanskii–Teissier inequality or the reverse Khovanskii–Teissier inequality, see [22].
@@
The symbol theorem for Lorentzian polynomials states that, if the symbol of a linear operator $T$ is a Lorentzian polynomial, then $T$ sends Lorentzian polynomials to Lorentzian polynomials [9, Theorem 3.2]. Theorem 4.7 can be used to derive a volume polynomial analogue of the symbol theorem for Lorentzian polynomials. [^8] The study of symbols of linear operators dates back to Gårding [14] and appears prominently in the work of Borcea and Brändén on the Pólya–Schur program for stable polynomials [7, 6].
@@
then, by [16, Lemma 2.1], it preserves the classes of irreducible cycles up to a rational multiple.
@@
    The upper truncation operators and the lower truncation operators preserve realizable volume polynomials over $k$ [16, Corollary 3.3].
@@
    The polarization operator $\Pi^{\uparrow}$ preserves realizable volume polynomials over $k$ [16, Proposition 4.1].
@@
    The normalization operator $\mathrm{N}$ preserves realizable volume polynomials over $k$ [16, Proposition 4.2].
@@
    For any nonnegative rational number $t$, the interlacing operator $1+tx_{i}\partial_{j}$ preserves realizable volume polynomials over $k$ [16, Proposition 4.3].
@@
    For any nonnegative rational number $t$, the symmetric exclusion process $\Phi^{i,j}_{t}$ preserves realizable multiaffine volume polynomials over $k$ [16, Proposition 4.4].
@@ 5. Realization problems for polymatroids
The following connection between algebraic matroids and the support of realizable volume polynomials is known [16, Proposition 5.4]. A polymatroid on $E$ is algebraic over $k$ if there are field extensions $k\subseteq\ell_{i}\subseteq\ell$ for $i\in E$ such that
@@
Proposition 5.5 implies that every minor of an algebraic polymatroid over $k$ is an algebraic polymatroid over $k$ [16, Section 5]. [^10] Definition 3.9, applied to the basis generating polynomials of polymatroids, define the notion of minor of polymatroids [2, Section 2.1]. In the classical case of matroids, this statement is typically deduced from a theorem of Lindström [29], who proved Piff’s conjecture that $M$ is algebraic over $k$ if $M$ is algebraic over an extension of $k$, see [32, Corollary 6.7.14]. Since the set of realizable volume polynomials over $k$ depends only on the characteristic of $k$ [16, Proposition 2.10], Proposition 5.5 gives the following version of Lindström’s theorem for polymatroids.
@@
Another consequence of Proposition 5.5 is that the intersection $M_{1}\wedge M_{2}$ of algebraic matroids over $k$ is an algebraic matroid over $k$ [16, Theorem 5.11]. This generalizes Piff’s theorem that the truncation of an algebraic matroid is algebraic [38, Section 11.3]. [^11] In [38, Section 11.3], Welsh proves the dual statement that the union $M_{1}\vee M_{2}$ of algebraic matroids over $k$ is an algebraic matroid over $k$. It is interesting to note that, as in Remark 4.3, the proof of the statement for $M_{1}\wedge M_{2}$ is substantially different from that for $M_{1}\vee M_{2}$.
Is the support of a realizable covolume polynomial over $k$ an algebraic polymatroid over $k$? This question extends the following long-standing open problem in matroid theory. For up-to-date discussions, see [5, 19].